OpenStudy (lilai3):

Who <333333333333s communnicaaatioooooooooon????????? How many phone numbers are possible for one area code if the first four numbers are 202-1, in that order, and the last three numbers are1-7-8 in any order?

OpenStudy (anonymous):

I think it is 3!

jimthompson5910 (jim_thompson5910):

There are 7 digits total. The first 4 are locked in at 202-1 So there are 7-4 = 3 digits left The last three numbers are 1-7-8 in any order and there are 3! = 3*2*1 = 6 ways to order them So there are 1*6 = 6 different phone numbers.

OpenStudy (lilai3):

Also for this question: Juan and Andrew are planning the schedule for a softball tournament. If there are 6 teams, how many different pairings could they make for the first tournament game? Is the answer to this one^^^^^ 6! which is 720 different pairings????? thank you got ur help!

OpenStudy (anonymous):

SimplyC(6,2)

jimthompson5910 (jim_thompson5910):

There are 6 teams to choose for the first slot (say the "home" slot) and 6-1 = 5 teams to choose for the second ("away/visitor") slot So there are 6*5 = 30 different ways to do this.

OpenStudy (lilai3):

oh so it's 30 instead for the answer to question number 2?

jimthompson5910 (jim_thompson5910):

If it's all being played on one field, then order doesn't matter, so you would divide by 2 to avoid the double counting.